OpenStudy (anonymous):

Can someone tell me the formula for newton's law of cooling and what doe each variable represent

OpenStudy (anonymous):

dT/dt=k(T-C) where T is the final temperature, C is the temperature of the surroundings and k is a constant

OpenStudy (anonymous):

It's T = A + Se^(-kt) where T = temperature at any instant A = temperature of surroundings (ambient temperature) S = initial temperature (I used S for starting) t= elapsed time and k is a constant determined by the conditions.